An object starts from rest and accelerated uniformly (constant acceleration). If it moves 3.0 m during the first second, then, during the first 5.00 seconds, it will move:

A) 15m B) 45m C) 75m D) 225m E) 1.0m

Equations to choose from:

x=xo+Vof +1/2at2

V=vo+at

V2=Vo2+2a(x-xo)
Answer #1

Here,

let the acceleration of the object is a

using the first equation

x=xo+Vof +1/2at2

3 = 0 + 0 * 1 + 0.50 * a * 1^2

a = 6 m/s^2

Now, for the distance travelled in the first 5 s

Distance = 0.50 * a * t^2

Distance = 0.50 * 6 * 5^2

Distance = 75 m

the distance traveled in first 5 s is C) 75 m